What is the pH of a buffer that consist of 0.20M NH3 and 0.10 M NH4Cl ? PKB of NH3= 4.75.
1
Expert's answer
2014-01-27T09:08:43-0500
pH = 14 - pKb + lg([NH3]/[NH4Cl])
pH = 14 - 4.75 + lg(0.2/0.1) = 9.55